# Shannon Capacity of Channels with Markov Insertions, Deletions and Substitutions

### v3

@@ II-E Information stability
III Functions of channels and their properties
$W^{\prime}(z|x)=\sum_{y\in f^{-1}(z)}W(y|x),$ (13)
where $f^{-1}(z)={\left\{{y:f(y)=z}\right\}}$ is the inverse image of $f$. Channel $W^{\prime}$ is degraded with respect to $W$, meaning that it can be represented as a pipeline of $W$ and a function $f$ at its output. It is well-known and can be straightforwardly checked that ${\mathbf{C}}(W^{\prime})\leq{\mathbf{C}}(W)$, and if both ${\mathbf{I}}(W)$ and ${\mathbf{I}}(W^{\prime})$ exist, then ${\mathbf{I}}(W^{\prime})\leq{\mathbf{I}}(W)$. Let $\Phi=\log_{2}\max_{z\in{\mathcal{Z}}}\left|{f^{-1}(z)}\right|$ be the maximum log-size of a preimage. Then, the following proposition holds.
Proposition 3 (Equation (4.3) in [5]).
For any input distribution $p\in{\mathbb{D}}_{{\mathcal{X}}}$:
${\mathbf{E}}_{p\circ W(x,y)}\bigg{[}\big{|}\mathbf{i}(W,p,x,y)-\mathbf{i}(W^{\prime},p,x,f(y))\big{|}\bigg{]}\leq\Phi.$ (14)
This proposition can be interpreted as follows. The process of forming the (random) output of channel $W$ can be seen as a pipeline of two steps. First, choose the value of $z$ based on probabilities $W^{\prime}(z|x)$. This is equivalent to channel $W^{\prime}$. Second, randomly choose one value among all (at most) $2^{\Phi}$ possible values of $y\in f^{-1}(y^{\prime})$. The maximum entropy of the second step is achieved when all preimages are equiprobable, namely, the entropy of the second step is at most $\Phi$.
For a channel sequence $U_{n}:{\mathcal{X}}^{n}\rightsquigarrow{\mathcal{Y}}_{n}$ and sequence of functions $f_{n}:{\mathcal{Y}}_{n}\to{\mathcal{Z}}_{n}$, we denote by $V=f(U)$ a channel sequence $V_{n}=f_{n}(U_{n}):{\mathcal{X}}^{n}\rightsquigarrow{\mathcal{Z}}_{n}$, $n\in{\mathbb{N}}$. Also, we denote by $\Phi_{n}=\log_{2}\max_{z\in{\mathcal{Z}}_{n}}\left|{f_{n}^{-1}(z)}\right|$.
Proposition 4 (Function with small preimage).
Proof.
First, note that ${\mathbf{E}}_{p_{n}\circ V_{n}}{\left[{\mathbf{i}(V_{n},p_{n},x,y)}\right]}={\mathbf{E}}_{p_{n}\circ U_{n}}{\left[{\mathbf{i}(V_{n},p_{n},x,f(y))}\right]}$. Using Proposition 3 and the fact that $V_{n}$ is degraded with respect to $U_{n}$, for any sequence of input distributions $p$:
00 0 $\leq\lim_{n\to\infty}\frac{{\mathbf{I}}(U_{n},p_{n})-{\mathbf{I}}(V_{n},p_{n})}{n}={\mathbf{E}}_{p_{n}\circ U_{n}}{\left[{\frac{\mathbf{i}(U_{n},p_{n},x,y)-\mathbf{i}(V_{n},p_{n},x,f_{n}(y))}{n}}\right]}$
$\leq\lim_{n\to\infty}{\mathbf{E}}_{p_{n}\circ U_{n}}{\left[{\frac{\left|{\mathbf{i}(U_{n},p_{n},x,y)-\mathbf{i}(V_{n},p_{n},x,f_{n}(y))}\right|}{n}}\right]}\leq\lim_{n\to\infty}\frac{\Phi_{n}}{n}=0,$ (15)
which for optimal $p$ for $U$ implies that ${\mathbf{I}}(U)={\mathbf{I}}(V)$, if ether ${\mathbf{I}}(U)$ or ${\mathbf{I}}(V)$ exists.
Equation (15) implies that for any $p_{n}\in{\mathbb{D}}_{{\mathcal{X}}^{n}}$,
$\lim_{n\to\infty}{\mathbf{E}}_{p_{n}\circ U_{n}}{\left[{\frac{\left|{\mathbf{i}(U_{n},p_{n},x,y)-\mathbf{i}(V_{n},p_{n},x,f_{n}(y))}\right|}{n}}\right]}=0.$
$\lim_{n\to\infty}{\Pr_{p_{n}\circ U_{n}}{\left\{{\frac{\left|{\mathbf{i}(U_{n},p_{n},x,y)-\mathbf{i}(V_{n},p_{n},x,f_{n}(y))}\right|}{n}>\delta}\right\}}}=0.$ (16)
$\lim_{n\to\infty}{\Pr_{p^{*}_{n}\circ U_{n}}{\left\{{\left|{\frac{\mathbf{i}(V_{n},p^{*}_{n},x,f_{n}(y))}{n}-{\mathbf{I}}(V)}\right|>\delta}\right\}}}\leq\lim_{n\to\infty}{\Pr_{p^{*}_{n}\circ U_{n}}{\left\{{\frac{\left|{\mathbf{i}(U_{n},p^{*}_{n},x,y)-\mathbf{i}(V_{n},p^{*}_{n},x,f_{n}(y))}\right|}{n}>\frac{\delta}{2}}\right\}}}$
$+\lim_{n\to\infty}{\Pr_{p^{*}_{n}\circ U_{n}}{\left\{{\left|{\frac{\mathbf{i}(U_{n},p^{*}_{n},x,y)}{n}-{\mathbf{I}}(U)}\right|>\frac{\delta}{2}}\right\}}}=0.$ (17)
To prove that ${\mathbf{C}}(U)={\mathbf{C}}(V)$, we will use the expression of c-capacity given by (9). Assume that ${\mathbf{C}}(U)=\alpha$ and ${\mathbf{C}}(V)\neq\alpha$. Then, ${\mathbf{C}}(V)<\alpha$ and
$\forall\varepsilon>0,\delta>0,\exists n^{*}_{0}:\forall n\geq n_{0},\exists p^{*}_{n}:~{}\Pr_{p^{*}\circ U_{n}(x,y)}{\left[{\frac{\mathbf{i}(U_{n},p^{*},x,y)}{n}>\alpha-\varepsilon}\right]}$ $>1-\delta.$ (18)
$\exists\varepsilon^{*}>0,\delta^{*}>0:\forall n_{0},\exists n^{*}\geq n_{0}:\forall p_{n},~{}\Pr_{p\circ U_{n^{*}}(x,y)}{\left[{\frac{\mathbf{i}(V_{n^{*}},p,x,f_{n^{*}}(y))}{n^{*}}\leq\alpha-\varepsilon^{*}}\right]}$ $\geq\delta^{*}.$ (19)
Since (18) holds for any $\varepsilon>0$, $\delta>0$, we can let $\varepsilon=\varepsilon^{*}/2$, $\delta=\delta^{*}/2$, where $\varepsilon^{*}$ and $\delta^{*}$ are from (19), and obtain simultaneously
$\exists\varepsilon^{*}>0,\delta^{*}>0:\begin{cases}\exists n^{*}_{0}:\forall n\geq n^{*}_{0},\exists p_{n}^{*}:~{}\Pr_{p_{n}^{*}\circ U_{n}(x,y)}{\left[{\mathbf{i}(U_{n},p_{n}^{*},x,y)/n>\alpha-\varepsilon^{*}/2}\right]}>1-\delta^{*}/2\\ \forall n_{0},\exists n^{*}\geq n_{0}:\forall p_{n^{*}},~{}\Pr_{p_{n^{*}}\circ U_{n^{*}}(x,y)}{\left[{\mathbf{i}(V_{n^{*}},p_{n^{*}},x,f_{n^{*}}(y))/{n^{*}}\leq\alpha-\varepsilon^{*}}\right]}\geq\delta^{*}\end{cases}$
  • for any $n_{0}$, pick up the specific value of $n^{*}\geq\max{\left\{{n_{0},n^{*}_{0}}\right\}}$, for which the second inequality holds;
  • since $n^{*}\geq n^{*}_{0}$, the first equality still holds for some specific selection of $p^{*}_{n^{*}}$. The second inequality will hold for $p_{n^{*}}^{*}$ as well, since it holds for any $p_{n^{*}}$.
The above procedure results in
$\exists\varepsilon^{*}>0,\delta^{*}>0:\forall n_{0},\exists n\geq n_{0},\exists p_{n}:\begin{cases}\Pr_{p_{n}\circ U_{n}(x,y)}{\left[{\mathbf{i}(U_{n},p_{n},x,y)/n>\alpha-\varepsilon^{*}/2}\right]}>1-\delta^{*}/2\\ \Pr_{p_{n}\circ U_{n}(x,y)}{\left[{\mathbf{i}(V_{n},p_{n},x,f_{n}(y))/{n}\leq\alpha-\varepsilon^{*}}\right]}\geq\delta^{*}\end{cases}.$
Note that both probabilities are defined over the same probability space. The sum of the two probabilities is greater than $21+\delta^{*}/$ 2, which means that
which contradicts (16). ∎
Thus, one can apply such function sequences to a channel sequence $W$, changing neither ${\mathbf{I}}(W)$, nor ${\mathbf{C}}(W)$, nor information stability property.
Proposition 5 (Semi-bijective function).
Consider channels $U:{\mathcal{X}}\rightsquigarrow{\mathcal{Y}}$ and $V:{\mathcal{X}}\rightsquigarrow{\mathcal{Z}}$, such that $V=f(U)$, $f:{\mathcal{Y}}\to{\mathcal{Z}}$. Consider partitions ${\mathcal{Y}}={\mathcal{A}}\sqcup{\mathcal{B}}$, ${\mathcal{Z}}={\mathcal{A}}^{\prime}\sqcup{\mathcal{B}}^{\prime}$, such that $f({\mathcal{A}})={\mathcal{A}}^{\prime}$ and $f$ is bijective between ${\mathcal{A}}$ and ${\mathcal{A}}^{\prime}$. Denote by $\beta(x)=\Pr_{U(y|x)}{\left[{y\in{\mathcal{B}}}\right]}$ the probability that the output of $U$ belongs to ${\mathcal{B}}$ given input $x\in{\mathcal{X}}$. If $\forall x\in{\mathcal{X}}:\beta(x)\leq\overline{\beta}$, then for any $p\in{\mathbb{D}}_{\mathcal{X}}$,
Proof.
Let $U^{\prime}:{\mathcal{X}}\rightsquigarrow{\mathcal{B}}$ be a channel equivalent to $U$ with an additional condition that the output belongs to ${\mathcal{B}}$, with the normalized transition probabilities
$U^{\prime}(y|x)=\frac{U(y|x)}{\beta(x)}.$
${\mathbf{E}}_{p\circ U(x,y)}{\left[{\mathbf{i}(U,p,x,y)}\right]}=\sum_{y\in{\mathcal{Y}}}\sum_{x\in{\mathcal{X}}^{n}}p(x)U(y|x)\mathbf{i}(U,p,x,y)$
$=\underbrace{\sum_{y\in{\mathcal{A}}}\sum_{x\in{\mathcal{X}}}p(x)U(y|x)\mathbf{i}(U,p,x,y)}_{a}+\underbrace{\sum_{y\in{\mathcal{B}}}\sum_{x\in{\mathcal{X}}}p(x)U(y|x)\mathbf{i}(U,p,x,y)}_{b}$
Let us analyze term $b$. First, introduce the modified distribution ${\widetilde{p}}(x)=p(x)\beta(x)/\Delta$, where $\Delta=\sum_{x}p(x)\beta(x)$. Note that $0\leq\Delta\leq\overline{\beta}$. Substituting $p(x)U(y|x)=\Delta{\widetilde{p}}(x)U^{\prime}(y|x)$ for $y\in{\mathcal{B}}$, one obtains
$b=\sum_{y\in{\mathcal{B}},x\in{\mathcal{X}}}p(x)U(y|x)\log_{2}\frac{U(y|x)}{\sum_{w}p(w)U(y|w)}=\Delta\sum_{y\in{\mathcal{B}},x\in{\mathcal{X}}}{\widetilde{p}}(x)U^{\prime}(y|x)\log_{2}\frac{\beta(x)U^{\prime}(y|x)}{\Delta\sum_{w}{\widetilde{p}}(w)U^{\prime}(y|w)}$
$=\Delta\cdot{\mathbf{I}}(U^{\prime},{\widetilde{p}})+\Delta\sum_{y\in{\mathcal{B}},x\in{\mathcal{X}}}{\widetilde{p}}(x)U^{\prime}(y|x)\log_{2}\frac{\beta(x)}{\Delta}=\Delta{\left({{\mathbf{I}}(U^{\prime},{\widetilde{p}})-\log_{2}\Delta}\right)}+\sum_{y\in{\mathcal{B}},x\in{\mathcal{X}}}p(x)U^{\prime}(y|x)\beta(x)\log_{2}\beta(x)$
The same can be done for the channel $V$:
${\mathbf{E}}_{p\circ V(x,y)}{\left[{\mathbf{i}(V,p,x,y)}\right]}=\sum_{y\in{\mathcal{A}}^{\prime}}\sum_{x\in{\mathcal{X}}}p(x)V(y|x)\mathbf{i}(V,p,x,y)+\underbrace{\sum_{y\in{\mathcal{B}}^{\prime}}\sum_{x\in{\mathcal{X}}}p(x)V(y|x)\mathbf{i}(V,p,x,y)}_{b^{\prime}}$
$=\sum_{y\in{\mathcal{A}}}\sum_{x\in{\mathcal{X}}}p(x)U(y|x)\mathbf{i}(U,p,x,y)+b^{\prime}=a+b^{\prime}$
$b^{\prime}=\sum_{z\in{\mathcal{B}}^{\prime},x\in{\mathcal{X}}}p(x)V(z|x)\mathbf{i}(V,p,x,z)=\Delta\sum_{z\in{\mathcal{B}}^{\prime},x\in{\mathcal{X}}}{\widetilde{p}}(x)V^{\prime}(z|x)\log_{2}\frac{\beta(x)V^{\prime}(z|x)}{\Delta\sum_{w}{\widetilde{p}}(w)V^{\prime}(z|w)}$
so (21) holds for $b^{\prime}$ as well. This implies
$\forall p\in{\mathbb{D}}_{\mathcal{X}}:{\mathbf{I}}(U,p)-{\mathbf{I}}(V,p)=b-b^{\prime}\leq\Delta\log_{2}\frac{|{\mathcal{X}}|}{\Delta}+\Delta\log_{2}\Delta+\frac{1}{e\ln 2}=\Delta\log_{2}|{\mathcal{X}}|+\frac{1}{e\ln 2}$
Since $\Delta\leq\overline{\beta}$, (20) holds. ∎
Proposition 6 (Function that does nothing almost surely).
Proof.
$0\leq{\mathbf{I}}(U)-{\mathbf{I}}(V)\leq\lim_{n\to\infty}\frac{{\mathbf{I}}(U_{n},p^{*}_{n})-{\mathbf{I}}(V_{n},p^{*}_{n})}{n}\leq\lim_{n\to\infty}{\left({\frac{\overline{\beta}_{n}n\log_{2}|{\mathcal{X}}|}{n}+\frac{1}{ne\ln 2}}\right)}=0,$
so ${\mathbf{I}}(U)={\mathbf{I}}(V)$.
Assume that ${\mathbf{I}}{\mathbf{S}}(U)$ is true. Note that, for any $p_{n}$,
$\lim_{n\to\infty}{\Pr_{p_{n}\circ U_{n}(x,y)}{\left\{{{\mathbf{i}}(U_{n},p_{n},x,y)\neq{\mathbf{i}}(V_{n},p_{n},x,f_{n}(y))}\right\}}}=\lim_{n\to\infty}{\Pr_{p_{n}\circ U_{n}(x,y)}{\left\{{y\in{\mathcal{B}}}\right\}}}\leq\lim_{n\to\infty}\overline{\beta}_{n}=0.$ (22)
$\lim_{n\to\infty}{\Pr_{p^{*}_{n}\circ V_{n}(x,z)}{\left\{{\left|{\frac{{\mathbf{i}}(V_{n},p^{*}_{n},x,z)}{n}-{\mathbf{I}}(V)}\right|>\delta}\right\}}}=\lim_{n\to\infty}{\Pr_{p^{*}_{n}\circ U_{n}(x,y)}{\left\{{\left|{\frac{{\mathbf{i}}(V_{n},p^{*}_{n},x,f_{n}(y))}{n}-{\mathbf{I}}(U)}\right|>\delta}\right\}}}$
$\leq\lim_{n\to\infty}\bigg{[}{\Pr_{p^{*}_{n}\circ U_{n}(x,y)}{\left\{{{\mathbf{i}}(V_{n},p^{*}_{n},x,f_{n}(y))\neq{\mathbf{i}}(U_{n},p^{*}_{n},x,y)}\right\}}}+{\Pr_{p^{*}_{n}\circ U_{n}(x,y)}{\left\{{\left|{\frac{{\mathbf{i}}(U_{n},p^{*}_{n},x,y)}{n}-{\mathbf{I}}(U)}\right|>\delta}\right\}}}\bigg{]},$
According to (9), ${\mathbf{C}}(U)={\mathbf{C}}(V)$ can be proved by analyzing the probability of $\mathbf{i}(V_{n},p_{n},x,y)$ being less than $\alpha n$ for some value of $\alpha$. This can be done as follows:
$\Pr_{p_{n}\circ V_{n}(x,y)}{\left[{\mathbf{i}(V_{n},p_{n},x,y)\leq\alpha n}\right]}=\Pr_{p_{n}\circ U_{n}(x,y)}{\left[{\mathbf{i}(V_{n},p_{n},x,f(y))\leq\alpha n}\right]}$
$\leq\Pr_{p_{n}\circ U_{n}(x,y)}{\left[{\mathbf{i}(V_{n},p_{n},x,f(y))\neq\mathbf{i}(U_{n},p_{n},x,y)\vee\mathbf{i}(V_{n},p_{n},x,f(y))=\mathbf{i}(U_{n},p_{n},x,y)\leq\alpha n}\right]}$
$\leq\Pr_{p_{n}\circ U_{n}(x,y)}{\left[{\mathbf{i}(V_{n},p_{n},x,f(y))\neq\mathbf{i}(U_{n},p_{n},x,y)\vee\mathbf{i}(U_{n},p_{n},x,y)\leq\alpha n}\right]}$
$\leq\Pr_{p_{n}\circ U_{n}(x,y)}{\left[{\mathbf{i}(V_{n},p_{n},x,f(y))\neq\mathbf{i}(U_{n},p_{n},x,y)}\right]}+\Pr_{p_{n}\circ U_{n}(x,y)}{\left[{\mathbf{i}(U_{n},p_{n},x,y)\leq\alpha n}\right]},$
where the first probability goes to zero according to (22). Thus,
$\lim_{n\to\infty}\min_{p_{n}}\Pr_{p_{n}\circ U_{n}(x,y)}{\left[{\mathbf{i}(U_{n},p_{n},x,y)\leq\alpha n}\right]}=0\implies\lim_{n\to\infty}\min_{p_{n}}\Pr_{p_{n}\circ V_{n}(x,y)}{\left[{\mathbf{i}(V_{n},p_{n},x,y)\leq\alpha n}\right]}=0,$
which means that the set of such $\alpha$’s for $U_{n}$ is a subset of such set for $V_{n}$, which implies ${\mathbf{C}}(V)\geq{\mathbf{C}}(U)$ and thus ${\mathbf{C}}(V)={\mathbf{C}}(U)$. ∎
If this proposition is combined with Proposition 4, the result is that one can apply any finite number of functions, which either have sufficiently small pre-image or sufficiently small probability of changing the output, without changing neither ${\mathbf{I}}$ nor ${\mathbf{C}}$ of the original channel sequence. The ${\mathbf{I}}{\mathbf{S}}$ property is also saved upon such transformations.
IV The Markov-IDS channel sequence
IV-A The IDS channel sequence
In [5] a general channel model, which represents any channel with independent insertions, deletions, and substitutions is considered. Consider a basic channel $W:{\mathcal{X}}\rightsquigarrow{\overline{\mathcal{Y}}}$, such that for each $x\in{\mathcal{X}}$ the expected length ${\mathbf{N}}(y)$ of the output sequence $y\in{\overline{\mathcal{Y}}}$ is finite:
@@ IV-A The IDS channel sequence
We will call such channel an IDS channel. The IDS channel sequence is $D=D^{\infty}$, where $D_{n}$ is channel $W^{n}$, followed by a gluing operator on its output:
@@
where $\mathbf{g}_{n}:{\overline{\mathcal{Y}}}^{n}\to{\overline{\mathcal{Y}}}$ glues together $n$ vectors over ${\mathcal{Y}}$. Such channel sequence $D$ is information-stable [5], so ${\mathbf{C}}(D)={\mathbf{I}}(D)$.
@@
Since all the terms of the summation are non-negative, we only need to prove that the sum is constrained. Denote by $p_{n}=\sum_{y\in{\mathcal{Y}}^{n}}w_{y}$ the probability that the length of output is $n$. From (23),
$\sum_{y\in{\overline{\mathcal{Y}}}}w_{y}{\mathbf{N}}(y)=\sum_{n=0}^{\infty}n\cdot\sum_{y\in{\mathcal{Y}}^{n}}w_{y}=\sum_{n=0}^{\infty}np_{n}\leq A.$
We obtained a constraint on $p_{n}$. Expression $\sum_{y\in{\mathcal{Y}}^{n}}-w_{y}\log_{2}w_{y}$ achieves its maximum (with fixed $p_{n}$) with uniform distribution within ${\mathcal{Y}}^{n}$, i.e., when $w_{y}=p_{n}/|{\mathcal{Y}}|^{n}$. Thus,
${\mathbf{H}}{\left[{w_{y}}\right]}\leq-\sum_{n=0}^{\infty}\sum_{y\in{\mathcal{Y}}^{n}}\frac{p_{n}}{|{\mathcal{Y}}|^{n}}\log_{2}\frac{p_{n}}{|{\mathcal{Y}}|^{n}}\leq{\mathbf{H}}{\left[{p_{n}}\right]}-A\cdot|{\mathcal{Y}}|.$
The second term is finite. The first term is the entropy of distribution $p_{n}$, corresponding to the length of the output of $W$, which has finite expectation $\leq A$. The maximum possible entropy of discrete distribution over ${\mathbb{N}}\cup{\left\{{0}\right\}}$ with given (finite) mean is equal to the entropy of the geometric distribution, which is finite. ∎
IV-B The Markov-IDS channel sequence
@@
The finite-state Markov chain (FSMC) channel [7] works as follows:
@@
$W_{n,1}[\sigma_{1}](y_{1}^{n}|x_{1}^{n})=W_{1}[\sigma_{1}](y_{1}|x_{1})\cdot\sum_{\sigma_{2}^{n}\in{\mathcal{S}}^{n-1}}\prod_{i=2}^{n}G_{\sigma_{i-1},\sigma_{i}}W_{1}[\sigma_{i}](y_{i}|x_{i}).$ (27)
Also, introduce channel $W_{n}[\sigma]$ which outputs $\mathbf{g}_{n}({\overline{y}}_{1}^{n})\in{\overline{\mathcal{Y}}}$, where ${\overline{y}}_{1}^{n}$ is the output of $W_{n,1}[\sigma]$:
@@
Assume that the initial state of Markov chain is chosen according to distribution $\rho\in{\mathbb{D}}_{s}$. Then, the corresponding FSMC channel sequence is denoted by $V_{*,1}[\rho]$, $V_{n,1}[\rho]:{\mathcal{X}}^{n}\rightsquigarrow{\overline{\mathcal{Y}}}^{n}$:
@@
The Markov-IDS channel sequence $V[\rho]$, $V_{n}[\rho]=\mathbf{g}_{n}(V_{n,1}[\rho]):{\mathcal{X}}^{n}\to{\overline{\mathcal{Y}}}$ is given by
@@
V Capacity theorem for Markov-IDS channel

### v5

@@ II-E Information Stability
III Markov-IDS Channels and the Main Results
III-A The IDS Channel Sequence
Consider an IDS channel $W:{\mathcal{X}}\rightsquigarrow{\overline{\mathcal{Y}}}$ as in [9], such that for each $x\in{\mathcal{X}}$ the expected length ${\mathbf{N}}(y)$ of the output sequence $y\in{\overline{\mathcal{Y}}}$ is finite:
@@ III-A The IDS Channel Sequence
The IDS channel $D_{n}$ is defined as channel $W^{n}$, followed by a gluing operator on its output (see definition (67)):
@@
where $\mathbf{g}_{n}:{\overline{\mathcal{Y}}}^{n}\to{\overline{\mathcal{Y}}}$ glues together $n$ vectors over ${\mathcal{Y}}$. The IDS channel sequence $D^{\infty}$ is information-stable [9], so ${\mathbf{C}}(D)={\mathbf{I}}(D)$.
@@
Since all the terms of the summation are non-negative, we only need to prove that the sum is constrained. Denote by $p_{n}=\sum_{y\in{\mathcal{Y}}^{n}}w_{y}$ the probability that the output length is $n$. From (22),
$\sum_{y\in{\overline{\mathcal{Y}}}}w_{y}\cdot{\mathbf{N}}(y)=\sum_{n=0}^{\infty}n\cdot\sum_{y\in{\mathcal{Y}}^{n}}w_{y}=\sum_{n=0}^{\infty}n\cdot p_{n}\leq A.$ (26)
Expression $\sum_{y\in{\mathcal{Y}}^{n}}(-w_{y}\log_{2}w_{y})$ achieves its maximum (with fixed $p_{n}$), when $w_{y}=p_{n}/|{\mathcal{Y}}|^{n}$. Thus,
${\mathbf{H}}{\left[{w_{y}}\right]}\leq-\sum_{n=0}^{\infty}|{\mathcal{Y}}|^{n}\cdot\frac{p_{n}}{|{\mathcal{Y}}|^{n}}\log_{2}\frac{p_{n}}{|{\mathcal{Y}}|^{n}}\leq{\mathbf{H}}{\left[{p_{n}}\right]}-A\cdot|{\mathcal{Y}}|.$ (27)
The second term is finite. The first term is the entropy of distribution $p_{n}$, corresponding to the length of the output of $W$, which has finite expectation $\leq A$. The maximum possible entropy of discrete distribution over ${\mathbb{N}}\cup{\left\{{0}\right\}}$ with a given (finite) mean is equal to the entropy of the geometric distribution, which is finite. ∎
III-B The Markov-IDS Channel Sequence
@@
The FSMC channel [29] works as follows:
@@
$W_{n,1}[\sigma_{1}](y_{1}^{n}|x_{1}^{n})=W_{1}[\sigma_{1}](y_{1}|x_{1})$
×$\times\sum_{\sigma_{2}^{n}\in{\mathcal{S}}^{n-1}}\prod_{i=2}^{n}G_{\sigma_{i-1},\sigma_{i}}W_{1}[\sigma_{i}](y_{i}|x_{i}).$ (28)
Introduce channel $W_{n}[\sigma]$ which outputs the glued output of $W_{n,1}[\sigma]$:
@@
Assume that the initial state of the Markov chain is chosen according to distribution $\rho\in{\mathbb{D}}_{s}$. Then, the corresponding FSMC channel sequence is denoted by $V_{n,1}[\rho]:{\mathcal{X}}^{n}\rightsquigarrow{\overline{\mathcal{Y}}}^{n}$:
@@
The Markov-IDS channel sequence $V_{n}[\rho]=\mathbf{g}_{n}(V_{n,1}[\rho])$ is given by gluing the output of the FSMC channel:
@@
III-C Markov Chain State Distribution Convergence